1. For a sample of 35 items from a population for which the standard deviation is σ= 20.5 , the sample mean is 458.0. At the 0.05 level of significance, test H0: μ= 450  versus H1: μ> 450  . Determine and interpret the p-value for the test.


1
Expert's answer
2022-02-24T06:00:12-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu=450"

"H_1:\mu>450"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z > 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{458-450}{20.5/\sqrt{35}}\approx2.308714"

Using the P-value approach:

The p-value is "p=P(Z>2.308714)=0.010480," and since "p=0.010480<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "450," at the "\alpha = 0.05" significance level.